### v1

@@ Appendix C Controlling for Conservation Policies
We include two key conservation policy variables in the analysis. First, we use the share of municipal territory designated as a sustainable-use protected area, specifically Environmental Protection Areas (Áreas de Proteção Ambiental, APAs), based on spatial overlap with units listed in Brazil’s National Registry of Conservation Units (CNUC). APAs are legal conservation units that allow for sustainable land use but restrict deforestation and land conversion. Second, we include an indicator for whether the municipality was designated as a Priority Municipality under the federal government’s conservation strategy launched in 2008. This policy targeted municipalities with the highest deforestation rates with intensified enforcement actions, including embargoes, fines, and credit restrictions.121212 As shown by Assunção and Rocha (2019), this prioritization was effective in significantly reducing deforestation. These conservation policy controls help ensure that other concurrent conservation initiatives do not confound the estimated effect of environmental enforcement on violence.
Table C.1 reports the estimated second-stage coefficients when we control for these additional conservation policies. Column (1) replicates the main specification while Column (2) includes the two conservation policy controls. The estimated effect of environmental enforcement on homicide remains negative and statistically significant, although the effect becomes slightly more negative (–0.748) and significance at the 10% level only. The first-stage F-statistic remains above conventional thresholds (14.25), mitigating concerns about weak instruments. These results reinforce the conclusion that the observed relationship between law enforcement and violence is not simply driven by the spatial targeting of conservation policy.
Table C.1: 2SLS — Conservation Policies Control

### v2

@@ Appendix E Controlling for Conservation Policies
We include two key conservation policy variables in the analysis. First, we use the share of municipal territory designated as a sustainable-use protected area, specifically Environmental Protection Areas (Áreas de Proteção Ambiental, APAs), based on spatial overlap with units listed in Brazil’s National Registry of Conservation Units (CNUC). APAs are legal conservation units that allow for sustainable land use but restrict deforestation and land conversion. Second, we include an indicator for whether the municipality was designated as a Priority Municipality under the federal government’s conservation strategy launched in 2008. This policy targeted municipalities with the highest deforestation rates with intensified enforcement actions, including embargoes, fines, and credit restrictions. [^16] As shown by Assunção and Rocha (2019), this prioritization was effective in significantly reducing deforestation. These conservation policy controls help ensure that other concurrent conservation initiatives do not confound the estimated effect of environmental enforcement on violence.
Table E.1 reports the estimated second-stage coefficients when we control for these additional conservation policies. Column (1) replicates the main specification while Column (2) includes the two conservation policy controls. The estimated effect of environmental enforcement on homicide remains negative and statistically significant, although the effect becomes slightly more negative (–0.748) and significance at the 10% level only. The first-stage F-statistic remains above conventional thresholds (14.25), mitigating concerns about weak instruments. These results reinforce the conclusion that the observed relationship between law enforcement and violence is not simply driven by the spatial targeting of conservation policy.
Table E.1: 2SLS — Conservation Policies Control